OpenStudy (anonymous):

Point R is located at (-5, 1), and S is located at (4, -6). Find the coordinates of the point that lies halfway between R and S. (5 points) Select one: a. (-2, -1) b. (-3, 5) c. (-0.5, -2.5) d. (-4.5, 3.5)

OpenStudy (anonymous):

@ganeshie8 @Nurali

OpenStudy (yttrium):

just get their average. \[(x,y) = (\frac{ x_{1}+x_{2} }{ 2 },\frac{ y_{1}+y_{2} }{ 2 })\]

OpenStudy (anonymous):

Do I have to plug in the points in the question

OpenStudy (yttrium):

Yup. Of course.

OpenStudy (anonymous):

i got (-3,-2)

OpenStudy (yttrium):

are you sure?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i did the correct calculations

OpenStudy (anonymous):

@ganeshie8 @Nurali

OpenStudy (yttrium):

(-5+4)/2 = ?? (-6+1)/2 = ??

OpenStudy (anonymous):

i got C then thanks a lot